# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ X Conclusion
In this paper, we established the asymptotic conversion theory between i.i.d. pure states in RTA for a compact Lie group. Our main result, namely the conversion rate formula, significantly extends a previous study on the $U(1)$ group to arbitrary compact Lie groups, thereby providing a unified framework for quantifying symmetry breaking across quantum systems with diverse symmetries. For example, it now covers cases involving multiple non-commutative conserved quantities. Notably, the derived conversion rate formula has led to a resolution of the long-standing Marvian–Spekkens conjecture, which had remained open for over a decade. We remark that the asymptotic conversion rate in RTA cannot be derived from the approach in general resource theory for reversible conversion [113] using the generalized Stein’s lemma [114, 115, 116, 117]. This is because this framework does not apply to RTA since the regularized relative entropy vanishes in RTA [13], as emphasized in [113, 116]. We also demonstrated that one can remove the constraints on the conversion rate arising from symmetry subgroups by supplementing the protocol with a sublinear number of additional resource states. We also explored the applicability of our approach to mixed-state conversion and derived upper bounds on asymmetry cost and distillable asymmetry. While our analysis in the mixed-state case is not comprehensive, it would pave the way for future investigations.
@@
Our result establishes the QGT as the complete asymmetry monotone within RTA, analogous to the role of entanglement entropy in entanglement theory. This finding provides an operational interpretation of the QGT, originally introduced as a metric on quantum state space and commonly employed as a topological indicator. Thus, it reveals a fundamental connection among symmetry breaking, the geometry of quantum states, and condensed matter physics from the viewpoint of quantum resource theories. Given the broad success of entanglement entropy as a theoretical tool for investigating quantum correlations [3, 4, 5], adopting the QGT for quantifying asymmetry breaking will offer a powerful approach to exploring deeper structures governed by symmetry and its breaking.
@@ Acknowledgements.
Appendix A Proof for the main theorem for any projective unitary representations
In the main text, we proved Eq. (8) for (non-projective) unitary representations of a compact Lie group. In Appendix A.1, we first review the result in [59] that relates the conversion rates for projective unitary representations to that for (non-projective) unitary representations and explain a formula for the conversion rate for projective unitary representations that are continuous but not differentiable. In Appendix A.2, we show that the QGTs at different points of a connected component of a Lie group are interrelated by a congruence transformation, which plays an essential role in proving Theorem 1 from Eq. (8). In Appendix A.3, we complete the proof for Eq. (6) in Theorem 1 and Eq. (8) for any projective unitary representations that are differentiable. Furthermore, Eq. (9) is proven Sec. A.4.
A.1 Differentiability of representation
@@
A.2 Relation among matrix inequalities between QGTs at different group elements for unitary representations
Let us prove the relation of the QGT at different points of a connected compact Lie group.
@@
Let $U$ be a (non-projective) unitary representation of a connected compact Lie group $G$. For each $g\in G$, there exists an invertible real matrix $V(g)$ independent of the representation $U$ such that $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}=V(g)^{\top}\mathcal{Q}^{\psi}V(g)$.
We remark the same argument for the congruence transformation of quantum Fisher information matrices can be found in [65].
@@
Let us first review several useful facts on a compact Lie group $G$ and its Lie algebra $\mathfrak{g}$. Let $GL(n,\mathbb{C})$ denote the general linear group. A closed subgroup of $GL(n,\mathbb{C})$ is referred to as a closed linear group. Any compact Lie group $G$ is isomorphic to a closed linear group (see, e.g., Corollary 4.22 in [118]). Through this identification, we can regard the exponential map from the Lie algebra $\mathfrak{g}$ to a compact Lie group $G$ as the matrix exponential. Now, let us fix an arbitrary basis $\{B_{\mu}\}_{\mu=1}^{\dim G}$ of $\mathfrak{g}$. For any $g\in G$, the set $\{g^{-1}B_{\mu}g\}_{\mu=1}^{\dim G}$ is also a basis of $\mathfrak{g}$. Therefore, we can introduce an invertible real matrix $V(g)$ by
$g^{-1}B_{\mu}g\eqqcolon\sum_{\nu=1}^{\dim G}V(g)_{\nu\mu}B_{\nu}.$ (124)
If we further assume that $G$ is a connected and compact Lie group, then the exponential map is subjective (see e.g., [118]), implying that there is $A_{g}\in\mathfrak{g}$ such that $g=\exp(\mathrm{i}A_{g})$. Therefore,
$g^{-1}B_{\mu}g=\exp(-\mathrm{i}A_{g})B_{\mu}\exp(\mathrm{i}A_{g})=\sum_{n=0}^{\infty}\frac{\mathrm{i}^{n}}{n!}[\cdots[[B_{\mu},\underbrace{A_{g}],A_{g}]\cdots,A_{g}]}_{n}.$ (125)
This means that the matrix $V(g)$ is determined by the Lie algebraic structure and is independent of the representation if $G$ is a connected compact Lie group.
Let us fix a local coordinate of $G$ in a neighborhood of the identity $e\in G$, which parametrizes the elements as $g(\bm{\lambda})$ such that $g(\bm{0})=e$. In this coordinate system, the QGT for $\psi$ is given by
@@
As an immediate corollary, we find that a matrix inequality between QGTs holds for any point of a connected compact Lie group $G$ if it holds at a point in $G$.
@@
Let $U$ and $U^{\prime}$ be a (non-projective) unitary representation of a connected compact Lie group $G$. If a matrix inequality $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}$ holds for a point $g\in G$, then it also holds for any $g\in G$.
@@
We can generalize the above corollary to any compact Lie group $G$. Let $G_{0}$ be the identity component of $G$. Since $G_{0}$ is a normal subgroup of $G$, we can introduce a quotient group $G/G_{0}$, which characterizes the disconnected group elements. In other words, the component of $G$ connected to $g\in G$, i.e., the set of group elements that are connected to $g\in G$ can be written as $G_{0}g$. Note that a compact Lie group $G$ has only a finite number of components (see, e.g., [64]). We denote the decomposition into the connected components by $G=\bigsqcup_{i=0}^{k}G_{i}$. We arbitrarily fix $g_{i}\in G_{i}$ for $i=0,\cdots,k$. Applying the argument in Corollary 19 for each connected component, we get
$\Longleftrightarrow\forall i=0,\cdots,k,\quad\forall g\in G_{i},\quad\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}$
$\Longleftrightarrow\forall i=0,\cdots,k,\quad\mathcal{Q}^{\mathcal{U}_{g_{i}}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g_{i}}(\phi)}.$ (131)
A.3 Conversion rates for projective unitary representations and (non-projective) unitary representations
@@
Let us fix a local coordinate $g(\bm{\lambda})$ in the neighborhood of the identity $e\in G$ such that $g(\bm{0})=e$. From the definition of $\tilde{U}$, we get
$\partial_{\mu}\tilde{U}(g(\bm{\lambda}))\biggl{|}_{\bm{\lambda}=\bm{0}}$ (133)
$=\frac{1}{\det(U(g(\bm{\lambda}))}\sum_{i=1}^{d}U(g(\bm{\lambda}))^{\otimes i-1}\otimes\partial_{\mu}U(g(\bm{\lambda}))\otimes U(g(\bm{\lambda}))^{\otimes d-i}\biggl{|}_{\bm{\lambda}=\bm{0}}+U(g(\bm{\lambda}))^{\otimes d}\partial_{\mu}\frac{1}{\det(U(g(\bm{\lambda}))}\biggl{|}_{\bm{\lambda}=\bm{0}}$ (134)
$=e^{-\mathrm{i}\theta}\sum_{i=1}^{d}I^{\otimes i-1}\otimes\partial_{\mu}U(g(\bm{\lambda}))\biggl{|}_{\bm{\lambda}=\bm{0}}\otimes I^{\otimes d-i}+c_{\mu}e^{\mathrm{i}\theta d}I^{\otimes d}\qquad\left(U(e)\eqqcolon e^{\mathrm{i}\theta}I,\,c_{\mu}\coloneqq\partial_{\mu}\frac{1}{\det(U(g(\bm{\lambda}))}\biggl{|}_{\bm{\lambda}=\bm{0}}\right).$ (135)
@@
$\frac{1}{d}\mathcal{Q}_{\mu\nu}^{\psi^{\otimes d}}$ $=\frac{1}{d}\sum_{i=1}^{d}\left(\braket{\partial_{\mu}U(g(\bm{\lambda}))\psi|\partial_{\nu}U(g(\bm{\lambda}))\psi}-\braket{\partial_{\mu}U(g(\bm{\lambda}))\psi|U(g(\bm{\lambda}))\psi}\braket{U(g(\bm{\lambda}))\psi|\partial_{\nu}U(g(\bm{\lambda}))\psi}\right)\biggl{|}_{\bm{\lambda}=\bm{0}}$
@@
We consider the case where $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$. Note that the case where $\mathrm{Sym}_{G}(\psi)\not\subset\mathrm{Sym}_{G}(\phi)$ is separately treated in Appendix A.4.
From Proposition 5 and Proposition 6, we have proven Eq. (8) for any (non-projective) unitary representations. For any projective unitary representations $U$ and $U^{\prime}$ on $\mathcal{H}$ and $\mathcal{H}^{\prime}$, we define unitary representations $\tilde{U}$ and $\tilde{U}^{\prime}$ of $G$ on $\mathcal{H}^{\otimes d}$ and $\mathcal{H}^{\prime\otimes d^{\prime}}$ by Eq. (122). From Eq. (123), we have
@@
By using the result in Appendix A.2, we prove Eq. (6) in Theorem 1.
@@
A.4 Symmetry subgroup and the conversion rate
@@
Appendix B Proofs of lemmas in the converse part
B.1 Proof of Lemma 3

### v4

@@ VIII Conclusion
In this paper, we established the asymptotic conversion theory between i.i.d. pure states in RTA for a compact Lie group. Our main result, namely the conversion rate formula, significantly extends a previous study on the $U(1)$ group to arbitrary compact Lie groups, thereby providing a unified framework for quantifying symmetry breaking across quantum systems with diverse symmetries. For example, it now covers cases involving multiple non-commutative conserved quantities. Notably, the derived conversion rate formula has led to a resolution of the long-standing Marvian–Spekkens conjecture, which had remained open for over a decade. We remark that the asymptotic conversion rate in RTA cannot be derived from the approach in general resource theory for reversible conversion [126] using the generalized Stein’s lemma [127, 128, 129, 130]. This is because this framework does not apply to RTA since the regularized relative entropy vanishes in RTA [13], as emphasized in [126, 129]. We also explored the applicability of our approach to mixed-state conversion and derived upper bounds on asymmetry cost and distillable asymmetry. While our analysis in the mixed-state case is not comprehensive, it would pave the way for future investigations.
@@
Our result establishes the QGT as the complete asymmetry measure within RTA, analogous to the role of entanglement entropy in entanglement theory. This finding provides an operational interpretation of the QGT, originally introduced as a metric on quantum state space and commonly employed as a topological indicator. Thus, it reveals a fundamental connection among symmetry breaking, the geometry of quantum states, and condensed matter physics from the viewpoint of quantum resource theories. Given the broad success of entanglement entropy as a theoretical tool for investigating quantum correlations [3, 4, 5], adopting the QGT for quantifying asymmetry breaking will offer a powerful approach to exploring deeper structures governed by symmetry and its breaking.
@@ Acknowledgements.
Appendix Contents
Appendix A $G$-covariant channel and conservation laws
In this section, we review the connection between $G$-covariance of a channel and conservation laws, established by the covariant Stinespring theorem [75, 14]. We denote by $\mathcal{L}(\mathcal{H})$ the set of all linear operators on a Hilbert space $\mathcal{H}$.
Theorem 10 (Theorem 25 in [14]).
[…]
$V\left(U(g)\otimes U_{E}(g)\right)=\left(U(g)\otimes U_{E}(g)\right)V,\quad\forall g\in G,$ (98)
such that
$\mathcal{E}(\cdot)=\mathrm{Tr}_{E}\left(V\left(\cdot\otimes\eta_{E}\right)V^{\dagger}\right).$ (99)
This theorem extends to the setting in which the input and output systems may differ. Although the proof is straightforward, we include it here for completeness:
Corollary 11.
Let $\mathcal{H}_{A},\mathcal{H}_{B}$ be finite-dimensional Hilbert spaces equipped with projective unitary representation $U_{A},U_{B}$ of a compact group $G$, and let $\mathcal{E}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ be a $G$-covariant channel. Then there exists finite-dimensional Hilbert spaces $\mathcal{H}_{E}$ and $\mathcal{H}_{E^{\prime}}$ with projective unitary representations $U_{E}$ and $U_{E^{\prime}}$ of $G$ such that $\mathcal{H}_{A}\otimes\mathcal{H}_{E}=\mathcal{H}_{B}\otimes\mathcal{H}_{E^{\prime}}$, a $G$-symmetric state $\eta_{E}$ on the system $E$, and a unitary operator $V$ on $\mathcal{H}_{A}\otimes\mathcal{H}_{E}$ satisfying
$V\left(U_{A}(g)\otimes U_{E}(g)\right)=\left(U_{B}(g)\otimes U_{E^{\prime}}(g)\right)V,\quad\forall g\in G,$ (100)
such that
$\mathcal{E}(\cdot)=\mathrm{Tr}_{E^{\prime}}\left(V\left(\cdot\otimes\eta_{E}\right)V^{\dagger}\right).$ (101)
Proof.
It is straightforward to check the channel in the form of Eq. (101) is $G$-covariant. Suppose that $\mathcal{E}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is $G$-covariant. Define a channel $\mathcal{F}:\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\to\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ by $\mathcal{F}(\cdot)\coloneqq\eta_{A}\otimes\left(\mathcal{E}\circ\mathrm{Tr}_{B}(\cdot)\right)$, where $\eta_{A}$ is a $G$-symmetric state. Since $\mathcal{F}$ is $G$-symmetric, from Theorem 10, there exist a finite-dimensional Hilbert space $\mathcal{H}_{F}$ carrying a unitary representation $U_{F}$ of $G$, a $G$-symmetric state $\eta_{F}$, and a unitary operator $V$ on $\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{F}$ satisfying
$V\left(U(g)_{A}\otimes U_{B}(g)\otimes U_{F}(g)\right)=\left(U(g)_{A}\otimes U_{B}(g)\otimes U_{F}(g)\right)V,\quad\forall g\in G,$ (102)
such that
$\mathcal{F}(\cdot)=\mathrm{Tr}_{F}\left(V(\cdot\otimes\eta_{F})V^{\dagger}\right).$ (103)
Fix any $G$-symmetric state $\eta_{B}$ on the system $B$. Since $\mathcal{E}(\cdot)=\mathrm{Tr}_{A}\circ\mathcal{F}(\cdot\otimes\eta_{B})$, we obtain
$\mathcal{E}(\cdot)=\mathrm{Tr}_{E^{\prime}}\left(V(\cdot\otimes\eta_{E})V^{\dagger}\right)$ (104)
where $E\coloneqq BF$, $E^{\prime}\coloneqq AF$, and $\eta_{E}\coloneqq\eta_{B}\otimes\eta_{F}$. The systems $E$ and $E^{\prime}$ carry projective unitary representations $U_{E}=U_{B}\otimes U_{F}$ and $U_{E^{\prime}}=U_{A}\otimes U_{F}$, which satisfies Eq. (100) due to Eq. (102). ∎
The condition in Eq. (100) means that the unitary evolution operator $V$ respects the symmetry described by $G$. When $G$ is a Lie group, by introducing the generators through differentiation of the projective representations (cf. Eq. (6)), this condition implies
$V\left(X_{A,\mu}+X_{E,\mu}\right)V^{\dagger}=X_{B,\mu}+X_{E^{\prime},\mu},\quad\forall\mu=1,\cdots,\dim G,$ (105)
i.e., additive conservation laws for the generators. In particular, if $G$ is connected, the additive conservation laws are equivalent to the intertwining condition Eq. (100). In this sense, a $G$-covariant channel characterizes the most general evolution of an open system subject to additive conservation laws, provided that no asymmetry is supplied from external systems.
Appendix B Proof for the main theorem for any projective unitary representations
In the main text, we provided the sketch of proof of Eq. (9) for (non-projective) unitary representations of a compact Lie group, while the technical proofs are provided in later sections, in Appendices C and D.
In this section, we show Theorem 1 from Eq. (9). In Appendix B.1, we first review the result in [64] that relates the conversion rates for projective unitary representations to that for (non-projective) unitary representations and explain a formula for the conversion rate for projective unitary representations that are continuous but not differentiable. In Appendix B.2, we show that the QGTs at different points of a Lie group are interrelated by a congruence transformation, which plays an essential role in proving Theorem 1 from Eq. (9). In Appendix B.3, we complete the proof for Eq. (8) in Theorem 1 and Eq. (9) for any projective unitary representations that are differentiable. Furthermore, Eq. (10) is proven Sec. B.4.
B.1 Differentiability of representation
@@
B.2 Relation among matrix inequalities between QGTs at different group elements for unitary representations
Let us prove the relation of the QGT at different points of a compact Lie group.
@@
Let $U$ be a unitary representation of a compact Lie group $G$. For each $g\in G$, there exists an invertible real matrix $V(g)$ independent of the representation $U$ such that $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}=V(g)^{\top}\mathcal{Q}^{\psi}V(g)$.
We remark the same argument for the congruence transformation of quantum Fisher information matrices can be found in [80] for connected compact Lie groups.
@@
Let us first review several useful facts on a compact Lie group $G$ and its Lie algebra $\mathfrak{g}$. Let $GL(n,\mathbb{C})$ denote the general linear group. A closed subgroup of $GL(n,\mathbb{C})$ is referred to as a closed linear group. Any compact Lie group $G$ is isomorphic to a closed linear group (see, e.g., Corollary 4.22 in [132]). For $g\in G$, we define a linear map $\mathrm{Ad}_{g}:\mathfrak{g}\to\mathfrak{g}$ by $\mathrm{Ad}_{g}(X)\coloneqq gXg^{-1}$. The map $\mathrm{Ad}_{g}$ is an invertible linear transformation of $\mathfrak{g}$ (see, e.g., Proposition 3.33 in [131]).
Now, let us fix an arbitrary basis $\{B_{\mu}\}_{\mu=1}^{\dim G}$ of $\mathfrak{g}$. For any $g\in G$, the set $\{g^{-1}B_{\mu}g\}_{\mu=1}^{\dim G}$ is also a basis of $\mathfrak{g}$. Therefore, we can introduce an invertible real matrix $V(g)$ by
$\mathrm{Ad}_{g^{-1}}(B_{\mu})\eqqcolon\sum_{\nu=1}^{\dim G}V(g)_{\nu\mu}B_{\nu}.$ (108)
Since $\mathrm{Ad}_{g}$ is a linear invertible transformation, the matrix $V(g)$ is invertible, is determined by the Lie algebraic structure, and is independent of the representation of $G$.
Let us fix a local coordinate of $G$ in a neighborhood of the identity $e\in G$, which parametrizes the elements as $g(\lambda)$ such that $g(0)=e$. In this coordinate system, the QGT for a pure state $|\psi=\ket{\psi}\bra{\psi}$ is given by
@@
As an immediate corollary, we find that a matrix inequality between QGTs holds for any point of a compact Lie group $G$ if it holds at a point in $G$.
@@
Let $U$ and $U^{\prime}$ be a unitary representation of a compact Lie group $G$. If a matrix inequality $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}$ holds for a point $g\in G$, then it also holds for any $g\in G$.
@@
In particular, in Eq. (113), we can always take $e\in G$ as a representative point. Thus, we get
$\forall g\in G,\quad\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}$ $\Longleftrightarrow\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}.$ (114)
B.3 Conversion rates for projective unitary representations and (non-projective) unitary representations
@@
Let us fix a local coordinate $g(\lambda)$ in the neighborhood of the identity $e\in G$ such that $g(0)=e$. From the definition of $\tilde{U}$, we get
$\partial_{\mu}\tilde{U}(g(\lambda))\biggl|_{\lambda=0}$ (116)
$=\frac{1}{\det(U(g(\lambda))}\sum_{i=1}^{d}U(g(\lambda))^{\otimes i-1}\otimes\partial_{\mu}U(g(\lambda))\otimes U(g(\lambda))^{\otimes d-i}\biggl|_{\lambda=0}+U(g(\lambda))^{\otimes d}\partial_{\mu}\frac{1}{\det(U(g(\lambda))}\biggl|_{\lambda=0}$ (117)
$=e^{-\mathrm{i}\theta}\sum_{i=1}^{d}I^{\otimes i-1}\otimes\partial_{\mu}U(g(\lambda))\biggl|_{\lambda=0}\otimes I^{\otimes d-i}+c_{\mu}e^{\mathrm{i}\theta d}I^{\otimes d}\qquad\left(U(e)\eqqcolon e^{\mathrm{i}\theta}I,\,c_{\mu}\coloneqq\partial_{\mu}\frac{1}{\det(U(g(\lambda))}\biggl|_{\lambda=0}\right).$ (118)
@@
$\frac{1}{d}\mathcal{Q}_{\mu\nu}^{\psi^{\otimes d}}$ $=\frac{1}{d}\sum_{i=1}^{d}\left(\braket{\partial_{\mu}U(g(\lambda))\psi|\partial_{\nu}U(g(\lambda))\psi}-\braket{\partial_{\mu}U(g(\lambda))\psi|U(g(\lambda))\psi}\braket{U(g(\lambda))\psi|\partial_{\nu}U(g(\lambda))\psi}\right)\biggl|_{\lambda=0}$
@@
We consider the case where $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$. Note that the case where $\mathrm{Sym}_{G}(\psi)\not\subset\mathrm{Sym}_{G}(\phi)$ is separately treated in Appendix B.4.
We now extend Eq. (9) to any unitary representations, given that Eq. (9) is proven for any (non-projective) unitary representations. For any projective unitary representations $U$ and $U^{\prime}$ on $\mathcal{H}$ and $\mathcal{H}^{\prime}$, we define unitary representations $\tilde{U}$ and $\tilde{U}^{\prime}$ of $G$ on $\mathcal{H}^{\otimes d}$ and $\mathcal{H}^{\prime\otimes d^{\prime}}$ by Eq. (106). From Eq. (107), we have
@@
By using the result in Appendix B.2, we prove Eq. (8) in Theorem 1.
@@
B.4 Symmetry subgroup and the conversion rate
@@
Appendix C Proof of the converse part
In this section, we provide the detailed proof of the converse part, i.e., Eq. (38). Section C.1 provides the discussion of the relation between the metric adjusted skew information and the QGT, and their asymptotic discontinuities, which provides complementary information to Sec. IV.1. Technical details of the proofs of the lemmas are provided in Sec. C.2.
C.1 Metric adjusted skew information and the QGT
We here first explain the reason why we adopted operator monotone functions that does not satisfy the symmetry condition in the main text. In the discussion of Petz’s monotone function $\braket{\cdot,\cdot}_{f,\rho}$, a symmetry condition $f(t)=tf(t^{-1})$ is often imposed to the monotone function [66], which is equivalent to $\braket{A,B}_{f,\rho}=\braket{B^{\dagger},A^{\dagger}}_{f,\rho}$ for any linear operators $A$ and $B$. However, as we shall see below, such a symmetric monotone metric does not contain the information of the anti-symmetric part of the QGT, which motivated us to introduce a family of operator monotone functions in $f_{q}$ in Eq. (25), which does not satisfy the symmetry condition if $q\neq 1/2$. Note that the special case for $q=1/2$ corresponds to $f_{\mathrm{SLD}}(x)=(1+x)/2$, which is associated with the symmetric logarithmic derivative (SLD) metric, playing a central role in quantum estimation theory.
Let $f^{(\mathrm{s})}$ be an operator monotone function satisfying $f^{(\mathrm{s})}(0)>0$ and the symmetry condition $f^{(\mathrm{s})}(t)=tf^{(\mathrm{s})}(t^{-1})$. From the definition of the monotone metric in Eq. (24), we find
$\frac{f^{(\mathrm{s})}(0)}{2}\|\mathrm{i}[\psi,O]\|_{f^{(\mathrm{s})},\psi}^{2}=\gamma^{\dagger}\left(\frac{1}{2}\left(\mathcal{Q}^{\psi}+\left(\mathcal{Q}^{\psi}\right)^{*}\right)\right)\gamma,$ (134)
for $O\coloneqq\gamma^{\dagger}X$. This means that it captures only the symmetric part of $\mathcal{Q}^{\psi}$.
This equation can be viewed as a special case of Eq. (134), where $\gamma$ is a real vector.
Since the metric adjusted skew information is quadratic in $H$, for the same reason as the QGT explained in the main text, it has an asymptotic discontinuity [13, 23, 24], which requires careful treatment when analyzing its asymptotic rates. In Ref. [69], the asymptotic behavior of metric adjusted skew information is studied, which enables to define valid asymmetry monotones obtained from asymptotic rates of the metric adjusted skew information. We here review the results in [69] since the proof of monotonicity of the QGT under asymptotic conversion are their variants.
For an arbitrary Hermitian operator $H$ we denote its i.i.d. extension by $H_{N}$, defined as $H_{N}\coloneqq\sum_{n=1}^{N}I^{\otimes n-1}\otimes H\otimes I^{\otimes N-n}$. For a sequence of states $\{\sigma_{N}\}_{N}$ such that $\limsup_{N\to\infty}T(\sigma_{N},\phi^{\otimes N})\leq\epsilon$ for $\epsilon>0$, and an operator monotone function $f^{(\mathrm{s})}$ satisfying $f^{(\mathrm{s})}(0)>0$ and the symmetry condition $f^{(\mathrm{s})}(t)=tf^{(\mathrm{s})}(t^{-1})$, Lemma 3 in [69] shows that under a certain regularity condition, if $\epsilon>0$ is sufficiently small, it holds
for all sufficiently large $N$, where $h$ is a real-valued function independent of $N$, satisfying $\lim_{\epsilon\to 0}h(\epsilon)=0$. The central limit theorem provides an intuitive explanation of this lemma. In the i.i.d. setting, the probability distribution of $\phi^{\otimes N}$ with respect to the eigenbasis of $H_{N}$ approaches to the normal distribution after properly redefining the random variable. It turns out that the left-hand side of Eq. (136) is approximated from below by the variance for a probability distribution modified due the the error between $\sigma_{N}$ and $\phi^{\otimes N}$. In general, modifying the distribution at values far from the mean can result in a considerable change in the variance. However, since the normal distribution has an exponentially small tail, it is impossible to significantly reduce the variance. In other words, the metric adjusted skew information approximately takes a local minimum around the i.i.d. pure states [69], expressed by Eq. (136). Note that although the definition of convergence in the state conversion in [69] is slightly different from ours, both yield the equivalent convertibility condition, as shown in Appendix C.2.2.
Lemma 17 (Restatement of Eq. (34)).
for all sufficiently large $N$.
Lemma 17 is proven by carefully modifying the proof of Lemma 3 in [69]. Since the proof is involved, the details are presented in Appendix C.2.3.
C.2 Proof of Lemmas in the converse part
C.2.1 Proof of Lemma 3